OpenStudy (anonymous):

**Trying to get the hang of this** help? Write the equation of the line that is perpendicular to the line y = x + 4 and passes through the point (-6, 3). A. y = x - 1 B. y = x + 3 C. y = -x - 1 D. y = -x + 3

jimthompson5910 (jim_thompson5910):

What is the slope of y = x + 4

OpenStudy (anonymous):

1 ?

jimthompson5910 (jim_thompson5910):

correct since y = x + 4 is the same as y = 1x + 4 it is in the form y = mx+b where m = 1, b = 4

jimthompson5910 (jim_thompson5910):

1 = 1/1 flip the fraction and flip the sign to get -1/1 which is just -1 so the perpendicular slope is -1

jimthompson5910 (jim_thompson5910):

we want this new line to pass through (-6,3) so x = -6 and y = 3 y = mx + b y = -1*x + b 3 = -1*6 + b 3 = -6 + b 3+6 = b 9 = b b = 9 Because m = -1 and b = 9, the final answer is y = -x + 9

jimthompson5910 (jim_thompson5910):

oh sorry my bad

jimthompson5910 (jim_thompson5910):

I made a typo

jimthompson5910 (jim_thompson5910):

y = mx + b y = -1*x + b 3 = -1*(-6) + b 3 = 6 + b 3 - 6 = b -3 = b b = -3 ------------------------------------------------------- m = -1 and b = -3 so the equation is y = -x - 3

OpenStudy (anonymous):

that makes so much more sense Thank you!!

jimthompson5910 (jim_thompson5910):

unfortunately I don't see y = -x - 3 listed, but that is definitely the answer

jimthompson5910 (jim_thompson5910):

so you'll have to bring it up with your teacher

OpenStudy (anonymous):

Gotcha !